OpenStudy (anonymous):

given parametric equation: x=cp y=c/p; P(cp,c/p) Tangent to the curve yp^2+x=2cp The perpendicular line from the origin meet the tangent at point T. Find the locus of T as p varies

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

i would if i could understand what this is asking

OpenStudy (anonymous):

lets make p a number, say 2

OpenStudy (anonymous):

actully at point P (x,y)=(cp,c/p) at point q (x,y)=(cq,c/q)

OpenStudy (anonymous):

oh what p is the parameter

OpenStudy (anonymous):

yaya...

OpenStudy (anonymous):

lets fix c at say 2 and see what this looks like

OpenStudy (anonymous):

ok now we have a picture. http://www.wolframalpha.com/input/?i=parametric+plot+%282t%2C2%2Ft%29

OpenStudy (anonymous):

x=cp -----1 y=c/p p=c/y-----2 subtitube 2 into 1 x=c(c/y) xy=c^2

OpenStudy (anonymous):

ok.. it is a hyperbola

OpenStudy (anonymous):

actually i want to know the meaning of "perpendicular line from the origin meet the tangent at point T" what i know is like this